# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v1

@@ 3 Problem Description
  • The arrival rate of parcels destinating for each outlet at each loading station is the same [^2] When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination.;
  • The time for robots dropping parcels into outlets is negligible;
  • Each outlet allows at most four robots in different directions releasing parcels simultaneously, and the drop-off operation by robots around one outlet does not affect each other;
@@
  • Workers are evenly distributed in the loading zone;
  • The transportation of parcels after being collected in the outlet is not considered in this research, which may affect the operations stage for RSSs with different physical structure; samples of single-tier and two-tier design are provided by Zou et al., (2021);
@@
(b) Topology of sample network
@@
To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation. Notably, we use the term ”aisle” instead of ”road” in the following content to clarify that each link in the network is single-lane. A sample grid network with compact aisle layout is depicted in Figure 3(a). It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles. Unit cells are evenly distributed along each aisle and serve for robot localization (in some robotic warehouses, distinct unit cells can be identified by QR codes). The unit cell is a square with a side length of $D$. We assume the length of a robot is not greater than $D$. It allows robot to spin turn in a grid. Outlets collect the parcels dropped by robots. Outlets are evenly distributed over the sorting zone and divided by aisles, and each outlet occupies one grid. This layout ensures that robots move along aisles can directly drop parcels into outlets without additional turning, while also achieving sufficient density. Figure 3(b) shows the topology of sample layout in Figure 3(a). Conflict nodes (blue dot) connect the vertical and horizontal aisles, and unloading nodes (red hexagon) are the points where robots release parcels. To improve traffic efficiency, all aisles are one-way and the directions of any two adjacent aisles are opposite, which is a common strategy in real. To ensure the connectivity of the map, each loading station in loading zone has independent entrance and exit, and the number of horizontal and vertical aisles should be even. Hence, the number of loading stations is no more than the number of aisles.
@@
Figure 4: Common types of conflicts: (a) a swapping conflict, (b) a vertex conflict, (c) a following conflict, and (d) a cycle conflict
@@
The sorting robots share the same kinematic parameters, featuring a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$. Additionally, these robots possess the capability of in-place rotation, characterized by an average angular velocity denoted as $\omega_{r}$, with no consideration for turning acceleration or deceleration. To ensure collision-free movement of robots, the robot management system needs to avoid conflicts between different robots during path finding. Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep; (d) Cycle conflict occurs among multiple robots, with each robot having a following conflict with another robot. Among these conflicts, both swapping conflicts and cycle conflicts necessitate a re-planning of paths. The occurrence of either can result in significant losses in system efficiency.
@@
Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage. In the site planning stage, decision maker concentrates on the division of zones for different functions in the warehouse. The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations. The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, decision maker should adaptively adjust the number of workers and robots to minimize operations costs, given real-time sorting demands. As a warehouse is usually rented for years, it is advisable to keep it unchanged during this stage. The decision variables of operations stage are the number of workers and robots in peak season $n_{w}^{H},n_{r}^{H}$ and in off-peak season $n_{w}^{L},n_{r}^{L}$, where $L$ and $H$ in superscript indicate that the sorting demand is at low and high levels, respectively; and $w$ and $r$ in subscript represent worker and robot, respectively.
@@
The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$. The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: With increased demand, a small sorting zone may lead to congestion, thereby increasing the service time for robots and raising the rental cost of robots in operations costs; whereas, expanding the size of RSS would escalate the facility costs. In sections 4-6, we will respectively present the traffic management of robots, conduct system performance analysis under the proposed traffic management framework, and validate the proposed algorithm and performance estimation formula. In section 7, a layout design optimization model is formulated to minimize the total costs of RSS with given demand.
@@ 4 Management of High-density Robot Traffic in RSS
Managing a large number of robots within a dense network poses significant challenges, especially when striving to balance system throughput performance with the computational load of algorithms under unpredictable traffic demands. In order to address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021). It enables the planning of uninterrupted scheduling for robot fleets. In this section, we start with a motivation example to illustrate why introducing rhythm into robot management is necessary. Subsequently, we summarize the key concepts inherited from the original RC. By leveraging the node-reservation idea in classical MAPF solver, we propose a new centralized control framework for sorting scenario, i.e., the RC-S scheme. RC-S presents high orderliness in practical performance, serving as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
4.1 Motivation example
In conveyor-based sorting systems, there is always a mainline through which goods move. When passing through the target outlets, the goods are pushed out by barriers to complete the sorting process. We aim to combine the stability of conveyor-based sorting systems with the flexibility and cost-effectiveness of RSS. Consider a conveyor system comprising two distinct types of conveyor segments: Type A, with a set length and flow speed, and Type B, featuring a turntable capable of altering directions as depicted in Figure 5(a). It is assumed that the transit time for goods on segment A is odd, whereas for conveyor B, this time is essentially negligible. Positioned along the mainline’s direction, the indices of segment B are sequentially labeled $1,2,3,$ and so forth. The orientation of each B segment coincides with the mainline when its index shares the same parity as the current time; if not, it shifts to align with the branch line. Consequently, by initiating the loading of goods onto the main line exclusively at even-numbered intervals, the goods will seamlessly proceed along the mainline. To divert them from the mainline, we could activate a barrier on segment B, blocking the goods for one unit of time when they reach it. By controlling the state of all the barriers, we can divert the goods to different branch lines. The production rhythm of such line is two units of time.
(a) Mainline-branch structure
(b) Network of conveyor segments
Figure 5: Conveyor system with two types of conveyor segment
The rhythmic manner remains applicable when applying the aforementioned control method to a four-line network in Figure 5(b), where adjacent lines always have opposite directions. Each line maintains consistent throughput when all the barriers are inactive. Activating a barrier at the appropriate time allows goods to switch to another line. However, this may conflict with the original rhythm of goods on the new line. We need a scheduling and reservation mechanism to avoid such conflicts.
@@ 4.2 Virtual platoon and cycle
If we use robots to replace the functionality of conveyors in the motivation example, then the problem becomes a MAPF issue. Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP). VPs represent the spatiotemporal slot that is generated in a rhythmic manner and is available for being occupied by vehicles. As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant distance between each other, resembling an endless series of train carriages connected by invisible ropes of equal length in each road. We discretize time based on the generation time of VPs at the entrance, and we refer to each segment as a cycle. Each entrance generates VPs at a predetermined pace, noted as the cycle length, which is equal to the time gap between the arrival time of two consecutive VPs at each node in the network. By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized. Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration.
Whenever a vehicle needs to pass through the network, the system reserves one unit of capacity from the corresponding VPs, and then releases it with an appropriate start cycle. Then that vehicle enters the network, following the movement of certain VPs within each route segments to reach its destination, and exits the network. Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity. Therefore, the trajectories are conflict-free without any blockage.
@@
RC is essentially a traffic control method, akin to a combination of globally coordinated traffic signal control and road network perimeter control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed technical information on the RC strategy, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
@@ 4.3 Concept of RC-S
In order to embody RC in the RSS, we regulate the behavior of robots, and propose the scheme of rhythmic control for sorting scenario, abbreviated as RC-S. RC-S can be regarded as a centralized MAPF algorithm. To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node. Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c},\mathcal{V}_{u},\mathcal{A})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$ and $\mathcal{A}$ stand for the set of conflict nodes, unloading nodes and links, respectively. To reduce traffic conflicts and deadlocks, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7. Furthermore, the behaviors of robots are limited to three actions: moving straight, dropping off parcel without stopping, and left turning. The joint movement inside a virtual platoon is not practical for robots. To avoid collision within a platoon, the capacity of each VP within a cycle is set to one robot. Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively. Under RC-S, robot enters the network through one aisle with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$. In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
@@
$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{ls}$, respectively:
@@
In addition, auxiliary road for vehicle turning in the original RC is removed, since the robot is capable to spin at the intersection grid for turning. The process of turning and transfer between VPs are demonstrated in Figure 8, which takes six consecutive phases. As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP, and accelerate to the speed of VP. Turning action entails four phases for both the original and subsequent VPs, totaling eight phases that are unavailable to other robots. However, the actual motion spans merely six consecutive phases, thus leading to a reduction in traffic capacity.
@@
Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process. The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed. In summary, incorporating expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
@@
A feasible path in RC-S is a sequence of VPs and their occupying cycles. The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two loading stations at the periphery of the network.
@@
When a robot performs a sortation task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{S}$, where $\mathcal{S}$ is the set of all loading stations. Let $\mathcal{O}$ represent the set of outlets. Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{S},j\in\mathcal{S},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$. To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{S}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{S},k\in\mathcal{O}}\mathcal{R}(j,k,i)$. In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{S}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
$s.t.$ $\sum_{i\in\mathcal{S}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,e}x_{i,k}^{r}\leq N_{\nu}^{e}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u},e\in\mathcal{E}$ (7)
$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{S},k\in\mathcal{O}$ (8)
$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{S}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{S}$ (9)
$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{S},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
@@
There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise. In the objective function, $c_{i,k}^{r}$ is the travel time on path $r\in\mathcal{R}_{i,k}$, and $\hat{c}_{i,k}$ is the penalty of delay. The objective function is to minimize the total costs in this cycle. Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{e}$ is the pre-set remaining capacity in node $\nu$ in cycle $e$, $N_{\nu}^{e}\leq 1$. Parameter $\delta_{i,k}^{r,\nu,e}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $e$ and zero otherwise. Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise. Constraint (9) is the conservation of flow constraint, ensuring a stable number of robots at each loading station.
@@
Algorithm 1 A heuristic method to implement RC-S
0:  Set the maximum working cycle $N_{c}$. Use a reservation table to record the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
0:  Set $t=1$;
0:  If the candidate list is empty, go to step 6; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$. Calculate the entry cycle of each feasible path in $\mathcal{R}(i,j,k)$, and choose the feasible path with the smallest entry cycle.
0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path; else, add the waiting time of the robot by $\tau_{c}$.
0:  Remove the robot from the candidate list and go to step 3.
0:  $t=t+1$. If $t>N_{c}$, end; else, add new VPs at the entrance of each aisle, labeled as unoccupied in the reservation table. Remove the VPs which have left the network. go to step 2.

### v2

@@ 3 Problem Description
  • The arrival rate of parcels destinating for each outlet at each loading station is the same (When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination. Additionally, outlets with extremely low and high arrival rates can be placed together to balance the flow);
  • The time for robots dropping parcels into outlets is negligible (To satisfy this assumption, robots need to release goods in advance based on its speed before reaching the target outlet, similar to the behavior of conveyor-based sorters. If it is difficult to achieve due to technical and cost issues, a 2-second drop-off time per delivery must be introduced (Zou et al.,, 2021));
  • The drop-off operation by robots around one outlet does not affect each other, i.e., each outlet allows at most four robots releasing parcels simultaneously;
@@
  • Loading stations are evenly distributed in the loading zone;
  • The transportation of parcels after being collected in the outlet is not considered in this research.
@@
(b) Graph representation of sample layout
@@
To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation. Notably, we use the term “aisle” instead of “road” in the following content to clarify that each road in the network is single-lane. A sample grid-based map in RSS is depicted in Figure 3(a). It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles. Each cell in the map has a side length of $D$. Robots locate themselves using QR codes at the center of each cell and can only move between adjacent cells. We assume that the length of sorting robots is less than $D$, allowing them to rotate within a cell. We use the term “outlet” to refer to the special cells designated for collecting parcels delivered by robots. Outlets are evenly distributed throughout the sorting zone, divided by aisles. This layout allows robots to deliver parcels to outlets on both sides of the aisle and ensures that the outlets are arranged with sufficient density. Each loading station in the loading zone is connected to a pair of dedicated entrance and exit, which are adjacent to each other. Therefore, the number of loading stations cannot exceed the the number of aisles. Figure 3(b) shows the graph representation of the sample layout in Figure 3(a). In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), unloading node (red hexagon), or entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes. Conflict nodes are at the intersections of aisles, unloading nodes are where robots deliver parcels, and entrance/exit nodes are at aisle ends.
@@
Figure 4: Common types of conflicts: (a) a swapping conflict, (b) a vertex conflict, (c) a following conflict, and (d) a deadlock conflict
@@
The sorting robots share the same kinematic parameters, with a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$. Robots possess the capability of in-place rotation, with angular velocity denoted as $\omega_{r}$. To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path finding. Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time. It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep. The latter is stationary or moving at a lower speed; (d) Deadlock conflict occurs among multiple robots, with each robot having a following conflict with another robot. Among these conflicts, both swapping conflicts and deadlock conflicts necessitate a re-planning of paths. The occurrence of either can result in significant losses in system efficiency.
Figure 5: Overview of research methodology
@@
Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage. In the site planning stage, warehouse manager concentrates on the division of zones for different functions in the warehouse. The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations. The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, warehouse manager should adaptively adjust the number of workers and robots to minimize operations costs, given expected sorting demands. Since a warehouse is usually rented for years, it is advisable to keep the zone division unchanged during its operation. The decision variables of operations stage are the number of workers and robots in operating period $\sigma$, noted as $n_{w}^{\sigma},n_{r}^{\sigma}$, respectively.
@@
The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$. The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: A small network leads to lower capacity, thereby increasing the waiting time of robots and raising the rental cost of robots in operations costs; whereas, an overly large network leads to increased site costs and longer travel distance. An overview of research methodology is shown in Figure 5. In sections 4-6, we present the traffic management of robots, conduct system efficiency analysis, and validate the performance and accuracy of estimation formula, respectively. In section 7, a layout design optimization model is formulated to minimize the total costs of RSS given the demand.
@@ 4 Management of High-density Robot Traffic in RSS
Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput performance with the computational load of algorithms under unpredictable traffic conditions. To address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021). It enables the planning of uninterrupted scheduling for robot fleets. In this section, we begin with the key concepts inherited from the original RC: the virtual platoon and the cycle. By leveraging the reservation mechanism, we propose a new centralized framework called Rhythmic Control in Sorting Scenario (RC-S). RC-S presents high orderliness and serves as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
@@ 4.1 Virtual platoon and cycle
Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP). VPs represent the spatiotemporal slot that is generated in a rhythmic manner and keep uniform linear motion. As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant safe distance between each other. Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration. We discretize time based on the generation time of VPs at the entrance, and each time interval is noted as a cycle. The duration of each time interval is the cycle length. By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized.
Whenever a vehicle needs to pass through a lane, it is mandated to follow the movement of one VP in that lane until it reaches its destination and exits the lane. Before entering the network, the vehicle must wait for an unoccupied VP at the entrance. It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity from the corresponding VPs in different segments of the route. The capacity is released when the vehicle leaves each VP. Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity. Consequently, the trajectories of vehicles are conflict-free.
@@
RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network perimeter control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed information on the RC framework, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
@@ 4.2 Concept of RC-S
In order to embody RC in the RSS, we regulate the behavior of robots, and propose the RC-S scheme, which can be regarded as a centralized MAPF solver. To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node. Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{e}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance/exit nodes and edges, respectively. To reduce intersections, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7. Based on the network structure, each VP can accommodate at most one AGV. Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively. Under RC-S, the behaviors of robots are limited to three actions: moving straight, dropping off parcel, and turning. Robots enter the network with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$.
In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
@@
$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
@@
RSS does not require auxiliary turning lanes since robots are capable to rotate within the intersection grid. The process of turning and transfer between VPs for robots are demonstrated in Figure 8. As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP and follows the movement of it. Turning action entails one cycle for both the original and subsequent VPs, leading to a reduction in traffic capacity.
@@
Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process. The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed $\omega_{r}$, ensuring that the robot can complete a 90-degree turn within one phase. In summary, using expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
@@
A feasible path in RC-S is a sequence of VPs and their occupying cycles. The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two active loading stations at the periphery of the network.
@@
Definition 1 defines the candidate spatio-temporal paths for robots under RC-S, where an active loading station indicates that a worker is currently engaged in tasks. When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{L}$, where $\mathcal{L}$ is the set of all loading stations. Let $\mathcal{O}$ represent the set of outlets. Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$. To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$. In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,l}x_{i,k}^{r}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},l\in\mathcal{C}$ (7)
$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{L},k\in\mathcal{O}$ (8)
$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{L}$ (9)
$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{L},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
@@
There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise. In the objective function, $c_{i,k}^{r}$ is the travel time on path $r\in\mathcal{R}_{i,k}$, and $\hat{c}_{i,k}$ is the penalty of delay. The objective function is to minimize the total costs in this cycle. Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{l}$ is the updated remaining capacity in node $\nu$ in cycle $l$, $N_{\nu}^{l}\leq 1$. Parameter $\delta_{i,k}^{r,\nu,l}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $l$ and zero otherwise. Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise. Constraint (9) is the conservation of flow constraint, ensuring a stable number of robots at each loading station.
@@
Algorithm 1 A heuristic method to implement RC-S
0:  Load the reservation table that records the status of VPs on the map within $N_{p}$ cycles. Set the maximum search range in each cycle as $N_{t}$. Calculate feasible path set $\mathcal{R}$.
0:  If the candidate list is empty, go to step 5; else, find the robot with the longest waiting time in the candidate list, obtain its origin $i$ and target outlet $k$. Calculate the entry cycle of each feasible path in $\mathcal{R}_{i,k}$, and choose the feasible path with the smallest entry cycle.
0:  If the entry cycle is less than $N_{t}$, reserve the VPs according to the feasible path, and assign the path to the robot; else, add the waiting time of the robot by $\tau_{c}$.
0:  Remove the robot from the candidate list and go to step 2.
0:  Record the newly generated VPs in the reservation table, labeled as unoccupied. Remove the VPs which have left the network. Output the assignment of feasible path.